# How to Create a Flat Ten or Eleven Dimensional Space-time in the Interior of an Asymptotically Flat Four Dimensional String Theory

### v2

@@ 1 Introduction
The rest of the paper is organized as follows. In section 2 we describe a background where we have a locally flat region with large internal $T^{6}$, but the string coupling in that region is also large. This is not quite the configuration we want since we want the string coupling to remain small or finite. We remedy this problem in section 3 by placing the black hole solution of section 2 in the background of a second black hole. There we show that it is not only possible to identify a large region where the local physics is described by type IIA string theory in ten dimensional flat space-time, but it is also possible to identify another large region where the local physics is described by M-theory in eleven dimensional flat space-time. In section 4 we show, how by taking a different background in the same theory, we can produce a large region where a local observer experiences type IIB string theory in ten dimensional flat space-time with moderate or small string coupling. We end in section 5 with some comments.
@@ 2 D0-brane black hole in type IIA on $T^{6}$
$\int d^{10}x\sqrt{-\det G}\left[e^{-2\phi}\left\{R_{G}+4G^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right\}-G^{\mu\rho}G^{\nu\sigma}F_{\mu\nu}F_{\rho\sigma}\right]\,,$ (2.1)
where $F$ is the field strength associated with the RR one form gauge field, $G_{\mu\nu}$ is the string metric and $\phi$ is the dilaton field. The D6-brane is magnetically charged under the RR one form gauge field, and the black hole solution carrying such charges is given by[5]:
@@
$e^{-2\phi}$ $=$ $e^{-2\phi_{0}}\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{\phi}}\,,\qquad r_{-}\leq r_{+}\,,$ (2.2)
@@
$x^{4},\cdots,x^{9}$ are coordinates along the compact directions, which we shall take to be periodic with period $2\pi$. The $e^{-\phi_{0}}$ factors were not present in the solution given in [5], but we have included it here using the scaling ‘symmetry’ $\phi\to\phi+\phi_{0}$, $F\to e^{-\phi_{0}}F$ under which the action scales by $e^{-2\phi_{0}}$ and the equations of motion remain unchanged. The parameters $Q$ and the mass $M$ of the black hole solution are related to $r_{\pm}$ via the relations[5]:
@@
We now compute the scalar $\phi$ and the radii $R$ of the compact directions in this region. They are given by
$e^{-2\phi}=e^{-2\phi_{0}}\left({r_{+}+b\over b+a}\right)^{3/2}\,,\qquad R=\left({b+a\over r_{+}+b}\right)^{1/4}\,.$ (2.7)
@@
From (2.7) we see that for $b,a<<r_{+}$, the sizes of the compact directions, measured in the string metric, is small and $e^{-2\phi}$ is large, which means that the string coupling is small. We shall now go to the T-dual description of this geometry, by making an $R\to\widetilde{R}=1/R$ transformation on all the circles. During this transformation the four dimensional dilaton $\phi_{4}$ remains the same, $\phi$ transforms to $\widetilde{\phi}$ accordingly, and the integrated flux of gauge field is reinterpreted as the D0-brane charge of the black hole. The near horizon moduli now take the form: [^3] Note that we could get this result directly by solving the equations of motion of low energy supergravity. The only reason for going via D6-branes is to lift the solution from [5].
$\widetilde{R}=R^{-1}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}}=e^{2\phi_{4}}\,\widetilde{R}^{6}=e^{2\phi_{0}}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (2.9)
We also note that in the new duality frame, the asymptotic radii of the circles remain 1 and the dilaton $\phi$ takes value $\phi_{0}$.
We see from (2.9) that for $b,a<<r_{+}$, the radii $\widetilde{R}$ of the compact directions, measured in the string metric, is large. However the string coupling $e^{\widetilde{\phi}}$ in this region is large. Hence we still have not reached our goal, which is to produce a region of space-time where the internal radii can be made (arbitrarily) large and the string coupling can be kept small or finite.
To address this issue, we note that if we are allowed to adjust $\phi_{0}$, then we could take $e^{2\phi_{0}}$ to be small to compensate for the large contribution from the last factor in the expression for $e^{2\widetilde{\phi}}$ in (2.9). However we would like to get a large ten dimensional space-time in a given four dimensional string theory for which all the asymptotic moduli are fixed, including the coupling $e^{\phi_{0}}$. So $\phi_{0}$ cannot be adjusted to achieve our goal. In the next section we shall describe a different black hole solution for which we can identify arbitrarily large locally flat regions of space-time where the string coupling can be made arbitrarily small, keeping all other moduli finite. Thus by placing the black hole considered in this section (the first black hole) in the background of the second black hole that we shall describe in the next section, we can vary the effective $\phi_{0}$ that enters the computation of this section by varying the position of the first black hole. This way we’ll be able to achieve our goal of creating a region where the space-time is approximately ten dimensional and have finite ten dimensional string coupling.
@@ 3 Black hole with winding and momentum charge
In this section we shall analyze a non-rotating black hole solution in type IIA string theory on $T^{6}$, carrying fundamental string winding and momentum charge along one of the circles. For simplicity we shall take the momentum and winding charges to be equal. We also take the asymptotic radii of the circles to be 1 and the asymptotic string coupling to be 1. The solution is given by a special case of the solutions given in [6]. [^4] Even though the solution in [6] was given for heterotic string theory, the result in type II string theory can be obtained easily from there by truncating the 28 dimensional charge vector to 12 dimensional charge vector and the $28\times 28$ matrix valued scalar moduli $M$ to a $12\times 12$ symmetric $SO(6,6)$ matrix. The solution given here is obtained from the one in [6] by setting to zero the rotation parameter $a$ and the angle $\alpha$. $\alpha=0$ corresponds to the choice of equal winding and momentum charges. Also $\Phi$ in [6] is our $2\phi$. The four dimensional string metric, the dilaton, the $12\times 12$ symmetric, $SO(6,6)$ matrix valued scalar $M$ and the non-vanishing component of the gauge fields are given by:
@@
The scaling that makes the metric almost locally flat and the invariant tensors almost zero is $m\to\lambda^{\prime}m$, $\rho\to\lambda^{\prime}\rho$, $t\to\lambda^{\prime}t$ for large $\lambda^{\prime}$. Again we shall not display the factors of $\lambda^{\prime}$ explicitly but keep in mind that we work with large $m$ and $\rho$ with $\rho/m>2$ fixed. Also we shall take $\beta$ large but fixed. In this limit the geometry becomes nearly flat and all the invariant tensors constructed from two or more derivatives of fields become nearly zero. The dilaton $\phi$ has the form:
$e^{-2\phi}\simeq m\,\rho^{-1}\cosh\beta\,.$ (3.2)
We now see that for large $\beta$ with fixed $\rho/m>2$, $e^{-2\phi}$ is large and hence $\phi$ becomes large and negative. Therefore in this region the string coupling $e^{\phi}$ is small. If we take the parameter $\lambda^{\prime}$ to be much larger than the parameter $\lambda$ used in the scaling in section 2, then the size of the black hole described in section 2, even though large, is small compared to the distance scale over which the fields change appreciably in the black hole background given in (3). Thus the black hole of section 2 will appear to be a point object from the perspective of the background 3 and it makes sense to place it at any point in the background (3). In particular we can place it at $\rho\sim m$. This configuration will evolve with time, but the time scale over which the fields will evolve will be of order $\lambda^{\prime}$ that can be made arbitrarily large. In that case $\phi_{0}$ appearing in (2.9) will be given by the $\phi$ in (3.2). Substituting this into (2.9) gives
$\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}}={\rho\over m\cosh\beta}\,\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (3.3)
Taking $a,b<<r_{+}$, $\rho\sim m$ and $\beta\sim\cosh^{-1}[\{r_{+}/(b+a)\}^{3/2}]>>1$, we can achieve our goal of getting large $\widetilde{R}$ and finite $\widetilde{\phi}$ i.e. a locally flat geometry where space-time is ten dimensional and the string coupling is finite.
@@
$R_{11}=e^{2\widetilde{\phi}/3}=\left({\rho\over m\cosh\beta}\right)^{1/3}\left({r_{+}+b\over b+a}\right)^{1/2}\,,\qquad R_{M}=e^{-\widetilde{\phi}/3}\,\widetilde{R}=\left({m\cosh\beta\over\rho}\right)^{1/6}\,.$ (3.4)
From this we see that by taking $\beta$ and $r_{+}/(a+b)$ large, and appropriately choosing their ratios, it is possible to make both $R_{11}$ and $R_{M}$ large. In that case we would create a large region where a local observer would experience eleven dimensional M-theory in almost flat space-time. In fact, for the same geometry, characterized by large values of $r_{+}/a$ and $\beta$, we can vary $b\equiv r-r_{+}$, labelling the region, to have either $\widetilde{R}$ large with $\widetilde{\phi}$ small or finite (IIA in flat space-time), or $R_{11}$ and $R_{M}$ large (M-theory in flat space-time).
@@ 4 Black hole with purely winding charge
In this section we shall describe how to construct a configuration that contains an arbitrarily large region where the local observer experiences type IIB string theory in flat ten dimensional space-time. For this we need to make use of a third black hole solution in type IIA string theory on $T^{6}$ that carries only winding charge along one of the circles. We shall take this to be the ninth circle. This solution can again be read out from the general solution given in [6]. [^5] The solution corresponds to setting $\alpha=\beta$ in [6]. The radius $R_{9}$ of the ninth circle appears inside the matrix $M$ as a $2\times 2$ bloc matrix of the form ${1\over 2}\pmatrix{R_{9}^{2}+R_{9}^{-2}&R_{9}^{2}-R_{9}^{-2}\cr R_{9}^{2}-R_{9}^{-2}&R_{9}^{2}+R_{9}^{-2}}\,.$ ) . Using this the four dimensional dilaton $\phi_{4}$ and the radius $R_{9}$ of the ninth direction associated with the solution in [6] for $\alpha=\beta$ takes the form:
@@
where we have introduced new parameter $\widehat{m}$ and new radial variable $\widehat{\rho}$ to distinguish them from the variables used in section 3. The other radii are kept fixed at 1. ${\widehat{\phi}_{0}}$ denotes the asymptotic value of $\phi_{4}$ for this solution. – this is an extra parameter that was set equal to zero in [6]. This is also the asymptotic value of the ten dimensional dilaton $\phi$ for this solution since $R_{9}$ approaches 1 in this limit and the other radii are also fixed at 1. On the other hand at a generic position labelled by $\widehat{\rho}$, the ten dimensional dilaton $\phi$ takes the form
$e^{2\phi}=e^{2\phi_{4}}\,R_{9}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1}\,.$ (4.3)
@@
$e^{2\phi}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,,\qquad R_{9}={1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1}\,.$ (4.5)
We now put the first black hole, described in section 2, at the position $\widehat{\rho}$ by taking $m,\rho>>\hat{m},\hat{\rho}>>r_{\pm},a,b>>1$. This amounts to replacing $e^{2\phi_{0}}$ in (2.9) by $e^{2\phi}$ given in (4.5). Also, the radius of the ninth circle will now be multiplied by $R_{9}$ given in (4.5) since this becomes the asymptotic value of this radius from the perspective of the solution (2.9). This gives the following solution:
@@
$e^{2\widetilde{\phi}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (4.6)
@@
$e^{2\overline{\phi}}=e^{2\widetilde{\phi}}\widetilde{R}_{9}^{-2}=m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\left({r_{+}+b\over b+a}\right)\,.$ (4.7)
In order to get type IIB string theory with large compactification radii and weak or moderate coupling, we need $\overline{R},\overline{R}_{9}>>1$ and $e^{\overline{\phi}}\stackrel{{\scriptstyle<}}{{\sim}}1$. If we choose $\rho\sim m$ and $\widehat{\rho}\sim\widehat{m}$, then this requires:

### v3

@@ 1 Introduction
The rest of the paper is organized as follows. In section 2 we describe a background where we have a locally flat region with large internal $T^{6}$, but the string coupling in that region is also large. This is not quite the configuration we want since we want the string coupling to remain small or finite. We remedy this problem in section 3 by placing the black hole solution of section 2 in the background of a second black hole. There we show that it is not only possible to identify a large region where the local physics is described by type IIA string theory in ten dimensional flat space-time, but it is also possible to identify another large region where the local physics is described by M-theory in eleven dimensional flat space-time. In section 4 we show, how by taking a different background in the same theory, we can produce a large region where a local observer experiences type IIB string theory in ten dimensional flat space-time with moderate or small string coupling. In section 5 we give an alternative construction by which we can achieve each of the three configurations – type IIA in ten dimensiona, M-theory in eleven dimensions and type IIB in ten dimensions by single black hole solutions. We end in section 6 with some comments.
@@ 2 D0-brane black hole in type IIA on $T^{6}$
$\int d^{10}x\sqrt{-\det G}\left[e^{-2\phi_{10}}\left\{R_{G}+4G^{\mu\nu}\partial_{\mu}\phi_{10}\partial_{\nu}\phi_{10}\right\}-G^{\mu\rho}G^{\nu\sigma}F_{\mu\nu}F_{\rho\sigma}\right]\,,$ (2.1)
where $F$ is the field strength associated with the RR one form gauge field, $G_{\mu\nu}$ is the string metric and $\phi_{10}$ is the dilaton field. The D6-brane is magnetically charged under the RR one form gauge field, and the black hole solution carrying such charges is given by[5]:
@@
$e^{-2\phi_{10}}$ $=$ $e^{-2\phi_{0}}\,\left[1-\left({r_{-}\over r}\right)\right]^{\gamma_{\phi}}\,,\qquad r_{-}\leq r_{+}\,,$ (2.2)
@@
$x^{4},\cdots,x^{9}$ are coordinates along the compact directions, which we shall take to be periodic with period $2\pi$. The $e^{-\phi_{0}}$ factors were not present in the solution given in [5], but we have included it here using the scaling ‘symmetry’ $\phi_{10}\to\phi_{10}+\phi_{0}$, $F\to e^{-\phi_{0}}F$ under which the action scales by $e^{-2\phi_{0}}$ and the equations of motion remain unchanged. The parameters $Q$ and the mass $M$ of the black hole solution are related to $r_{\pm}$ via the relations[5]:
@@
We now compute the scalar $\phi_{10}$ and the radii $R$ of the compact directions in this region. They are given by
$e^{-2\phi_{10}}=e^{-2\phi_{0}}\left({r_{+}+b\over b+a}\right)^{3/2}\,,\qquad R=\left({b+a\over r_{+}+b}\right)^{1/4}\,.$ (2.7)
@@
From (2.7) we see that for $b,a<<r_{+}$, the sizes of the compact directions, measured in the string metric, is small and $e^{-2\phi_{10}}$ is large, which means that the string coupling is small. We shall now go to the T-dual description of this geometry, by making an $R\to\widetilde{R}=1/R$ transformation on all the circles. During this transformation the four dimensional dilaton $\phi_{4}$ remains the same, $\phi_{10}$ transforms to $\widetilde{\phi}_{10}$ accordingly, and the integrated flux of gauge field is reinterpreted as the D0-brane charge of the black hole. The near horizon moduli now take the form: [^3] Note that we could get this result directly by solving the equations of motion of low energy supergravity. The only reason for going via D6-branes is to lift the solution from [5].
$\widetilde{R}=R^{-1}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}_{10}}=e^{2\phi_{4}}\,\widetilde{R}^{6}=e^{2\phi_{0}}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (2.9)
We also note that in the new duality frame, the asymptotic radii of the circles remain 1 and the dilaton $\widetilde{\phi}_{10}$ takes value $\phi_{0}$.
We see from (2.9) that for $b,a<<r_{+}$, the radii $\widetilde{R}$ of the compact directions, measured in the string metric, is large. However the string coupling $e^{\widetilde{\phi}_{10}}$ in this region is large. Hence we still have not reached our goal, which is to produce a region of space-time where the internal radii can be made (arbitrarily) large and the string coupling can be kept small or finite.
To address this issue, we note that if we are allowed to adjust $\phi_{0}$, then we could take $e^{2\phi_{0}}$ to be small to compensate for the large contribution from the last factor in the expression for $e^{2\widetilde{\phi}_{10}}$ in (2.9). However we would like to get a large ten dimensional space-time in a given four dimensional string theory for which all the asymptotic moduli are fixed, including the coupling $e^{\phi_{0}}$. So $\phi_{0}$ cannot be adjusted to achieve our goal. In the next section we shall describe a different black hole solution for which we can identify arbitrarily large locally flat regions of space-time where the string coupling can be made arbitrarily small, keeping all other moduli finite. Thus by placing the black hole considered in this section (the first black hole) in the background of the second black hole that we shall describe in the next section, we can vary the effective $\phi_{0}$ that enters the computation of this section by varying the position of the first black hole. This way we’ll be able to achieve our goal of creating a region where the space-time is approximately ten dimensional and have finite ten dimensional string coupling.
@@ 3 Black hole with winding and momentum charge
In this section we shall analyze a non-rotating black hole solution in type IIA string theory on $T^{6}$, carrying fundamental string winding and momentum charge along one of the circles. For simplicity we shall take the momentum and winding charges to be equal. We also take the asymptotic radii of the circles to be 1 and the asymptotic string coupling to be 1. The solution is obtained from the solutions given in [6] and reviewed in appendix A by setting $\alpha=0$ and taking the unit vector $\vec{p}$ to be along the circle carrying the winding and momentum charges. The four dimensional string metric $ds_{4}^{2}$, the ten dimensional dilaton $\phi_{10}$, the four dimensional dilaton $\phi_{4}$, the $12\times 12$ symmetric, $SO(6,6)$ matrix valued scalar $M$ and the non-vanishing component of the gauge fields are given by:
@@
The scaling that makes the metric almost locally flat and the invariant tensors almost zero is $m\to\lambda^{\prime}m$, $\rho\to\lambda^{\prime}\rho$, $t\to\lambda^{\prime}t$ for large $\lambda^{\prime}$. Again we shall not display the factors of $\lambda^{\prime}$ explicitly but keep in mind that we work with large $m$ and $\rho$ with $\rho/m>2$ fixed. Also we shall take $\beta$ large but fixed. In this limit the geometry becomes nearly flat and all the invariant tensors constructed from two or more derivatives of fields become nearly zero. The dilaton $\phi_{10}$ has the form:
$e^{-2\phi_{10}}\simeq m\,\rho^{-1}\cosh\beta\,.$ (3.2)
We now see that for large $\beta$ with fixed $\rho/m>2$, $e^{-2\phi_{10}}$ is large and hence $\phi_{10}$ becomes large and negative. Therefore in this region the string coupling $e^{\phi_{10}}$ is small. If we take the parameter $\lambda^{\prime}$ to be much larger than the parameter $\lambda$ used in the scaling in section 2, then the size of the black hole described in section 2, even though large, is small compared to the distance scale over which the fields change appreciably in the black hole background given in (3). Thus the black hole of section 2 will appear to be a point object from the perspective of the background 3 and it makes sense to place it at any point in the background (3). In particular we can place it at $\rho\sim m$. This configuration will evolve with time, but the time scale over which the fields will evolve will be of order $\lambda^{\prime}$ that can be made arbitrarily large. In that case $\phi_{0}$ appearing in (2.9) will be given by the $\phi_{10}$ in (3.2). Substituting this into (2.9) gives
$\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad e^{2\widetilde{\phi}_{10}}={\rho\over m\cosh\beta}\,\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (3.3)
Taking $a,b<<r_{+}$, $\rho\sim m$ and $\beta\sim\cosh^{-1}[\{r_{+}/(b+a)\}^{3/2}]>>1$, we can achieve our goal of getting large $\widetilde{R}$ and finite $\widetilde{\phi}_{10}$ i.e. a locally flat geometry where space-time is ten dimensional and the string coupling is finite.
@@
$R_{11}=e^{2\widetilde{\phi}_{10}/3}=\left({\rho\over m\cosh\beta}\right)^{1/3}\left({r_{+}+b\over b+a}\right)^{1/2}\,,\qquad R_{M}=e^{-\widetilde{\phi}_{10}/3}\,\widetilde{R}=\left({m\cosh\beta\over\rho}\right)^{1/6}\,.$ (3.4)
From this we see that by taking $\beta$ and $r_{+}/(a+b)$ large, and appropriately choosing their ratios, it is possible to make both $R_{11}$ and $R_{M}$ large. In that case we would create a large region where a local observer would experience eleven dimensional M-theory in almost flat space-time. In fact, for the same geometry, characterized by large values of $r_{+}/a$ and $\beta$, we can vary $b\equiv r-r_{+}$, labelling the region, to have either $\widetilde{R}$ large with $\widetilde{\phi}_{10}$ small or finite (IIA in flat space-time), or $R_{11}$ and $R_{M}$ large (M-theory in flat space-time).
@@ 4 Black hole with purely winding charge
In this section we shall describe how to construct a configuration that contains an arbitrarily large region where the local observer experiences type IIB string theory in flat ten dimensional space-time. For this we need to make use of a third black hole solution in type IIA string theory on $T^{6}$ that carries only winding charge along one of the circles. We shall take this to be the ninth circle. This solution can again be read out from the general solution given in appendix A by setting $\alpha=\beta$ and taking the vectors $\vec{n}$ and $\vec{p}$ to be directed along the same compact direction $x^{9}$. Using (A)-(A.3), the four dimensional dilaton $\phi_{4}$ and the radius $R_{9}$ of the ninth direction takes the form:
@@
where we have introduced new parameter $\widehat{m}$ and new radial variable $\widehat{\rho}$ to distinguish them from the variables used in section 3. The other radii are kept fixed at 1. ${\widehat{\phi}_{0}}$ denotes the asymptotic value of $\phi_{4}$ for this solution. This is also the asymptotic value of the ten dimensional dilaton $\phi_{10}$ for this solution since $R_{9}$ approaches 1 in this limit and the other radii are also fixed at 1. On the other hand at a generic position labelled by $\widehat{\rho}$, the ten dimensional dilaton $\phi_{10}$ takes the form
$e^{2\phi_{10}}=e^{2\phi_{4}}\,R_{9}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1}\,.$ (4.3)
The horizon of this solution is at $\widehat{\rho}=2\widehat{m}$. Hence we shall stay in the region $\widehat{\rho}>2\widehat{m}$.
@@
$e^{2\phi_{10}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,,\qquad R_{9}={1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1}\,.$ (4.5)
We now put the first black hole, described in section 2, at the position $\widehat{\rho}$ by taking $m,\rho>>\hat{m},\hat{\rho}>>r_{\pm},a,b>>1$. This amounts to replacing $e^{2\phi_{0}}$ in (2.9) by $e^{2\phi_{10}}$ given in (4.5). Also, the radius of the ninth circle will now be multiplied by $R_{9}$ given in (4.5) since this becomes the asymptotic value of this radius from the perspective of the solution (2.9). This gives the following solution:
@@
$e^{2\widetilde{\phi}_{10}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (4.6)
@@
$e^{2\overline{\phi}_{10}}=e^{2\widetilde{\phi}_{10}}\widetilde{R}_{9}^{-2}=m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\left({r_{+}+b\over b+a}\right)\,.$ (4.7)
In order to get type IIB string theory with large compactification radii and weak or moderate coupling, we need $\overline{R},\overline{R}_{9}>>1$ and $e^{\overline{\phi}_{10}}\stackrel{{\scriptstyle<}}{{\sim}}1$. If we choose $\rho\sim m$ and $\widehat{\rho}\sim\widehat{m}$, then this requires:
@@
5 Achieving the goals with single black holes
In the constructions described above we made use of multiple black holes to produce arbitrarily large regions where a local observer experiences ten dimensional type IIA or type IIB string theories or eleven dimensional M-theory. It is natural to ask if this can be achieved using a single black hole. In this section we can describe how this can be done.
For this analysis we shall make use of black holes carrying charges representing momenta along the compact directions. We shall first describe the construction in type IIA string theory on $T^{6}$. This can again be obtained as a special case of the solution given in appendix A where we set $\alpha=-\beta$ and take the vectors $\vec{p}=\vec{n}$ to be some generic six dimensional unit vector. Then the four dimensional metric $ds_{4}^{2}$, the four dimensional dilaton $\phi_{4}$ and the internal six dimensional metric $ds_{6}^{2}$ describing the moduli fields have the following functional dependence on the radial variable $\rho$::
$\displaystyle ds_{4}^{2}$ $=$ $-{\rho-2m\over\rho+2m\sinh^{2}\alpha}dt^{2}+{\rho\over\rho-2m}d\rho^{2}+\rho^{2}d\Omega_{2}^{2}\,,$
$e^{2\phi_{4}}$ $=$ $e^{2\phi_{0}}\,(1+2m\rho^{-1}\sinh^{2}\alpha)^{-1/2},$
[…]
where $\phi_{0}$ is the asymptotic value of the dilaton $\phi_{4}$. The horizon of the solution is at $\rho=2m$ and the momentum charge vector carried by the black hole is proportional to $m\,\sinh^{2}\alpha\,\vec{n}$. As before, we shall take the large $m$ limit at fixed $\rho/m$ so as to smoothen the geometry and make it locally flat. From (5) we can also compute the ten dimensional dilaton
$e^{2\phi_{10}}=e^{2\phi_{4}}\,R=e^{2\phi_{0}}\,.$ (5.2)
Thus the ten dimensional dilaton is constant. This shows that this is a purely gravitational solution and hence can be easily lifted to any theory, e.g. M-theory on $T^{7}$ or type IIB on $T^{6}$.
By taking $\rho\sim m$ and $\alpha$ large, we can take $R$ to be large. In this case, apparently (5) indicates that only one of the directions $\vec{n}$ of the torus becomes large and the orthogonal directions remain finite. For example, if each $y^{i}$ is periodic with period one, and if $\vec{n}\propto(1,1,\cdots 1)$, then the cycle that shifts $y^{1}$ by $2\pi$ and $y^{2}$ by $-2\pi$ will have finite size since it will not be affected by the term proportional to $R^{2}$ in the metric. However by choosing $\vec{n}$ appropriately we can ensure that all closed cycles on the torus have large size. We shall illustrate this by an example below. Let us take
$\vec{n}=(1,c,c^{2},\cdots c^{d-1})/\sqrt{1+c^{2}+\cdots+c^{2d-2}}$ (5.3)
[…]
$2\pi\left[m_{1}^{2}+\cdots+m_{d}^{2}+(R^{2}-1)(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})^{2}/(1+c^{2}+\cdots+c^{2d-2})\right]^{1/2}\,.$ (5.4)
For generic $m_{i}$ the cycle will have size of order $R$. We could avoid it by choosing the $m_{i}$’s so that $(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})$ is small. This can be done in two ways: either set some of the $m_{i}$’s to zero or choose the $m_{i}$’s such that the different terms in the sum cancel. In the first approach the smallest value is achieved by setting $m_{i}=0$ for $i\neq d-1$ and $m_{d}=1$. This gives a cycle size of order $2\pi Rc^{(d-1)}\sim R^{1/d}$ for $c\sim R^{-1/d}$. Thus we have a large size for large $R$. On the other hand in order to have cancellation between different terms in $(m_{1}+m_{2}c+\cdots m_{d}c^{d-1})$, some of the $m_{i}$’s must be of order $c^{-1}$. Then the $R$ independent terms in (5.4) will have large contribution and we still will have a minimum cycle size of order $R^{1/d}$.
One can also see this by analyzing the spectrum of momentum carrying states. If $\vec{k}$ denotes the quantized momentum along $T^{d}$ taking value on a lattice of integers, then for the metric given in (5), its contribution to the mass2 in four space-time dimensions is given by:
$m^{2}=G^{ij}k_{i}k_{j}=\vec{k}^{2}+(R^{-2}-1)(\vec{n}.\vec{k})^{2}\,.$ (5.5)
Let us for definiteness take
$R=N^{d},\qquad c=N^{-1}\,,$ (5.6)
for some large integer $N$. This can be achieved by taking the momenta along the $d$ circles to be some multiple of $(1,N,N^{2},\cdots N^{d-1})$. We now introduce the following basis states in the momentum lattice:
$\vec{e}_{1}=(1,0,\cdots 0),\quad\vec{e}_{2}=(N,1,0,\cdots 0),\quad\vec{e}_{3}=(N^{2},N,1,0,\cdots 0),\quad\cdots,\ \vec{e}_{d}=(N^{d-1},N^{d-2},\cdots 1)\,.$ (5.7)
We also have from (5.3)
$\vec{n}=(1+N^{2}+\cdots N^{2d-2})^{-1/2}(N^{d-1},N^{d-2},\cdots 1)\,.$ (5.8)
Then a general momentum vector may be expressed as
$\vec{k}=\sum_{i=1}^{d}q_{i}\vec{e}_{i},\qquad q_{i}\in{\hbox{ Z\kern-4.55244ptZ}}\,.$ (5.9)
Substituting this into (5.5), we get
$m^{2}=N^{-2}\sum_{i=1}^{d}q_{i}^{2}+{\cal O}(N^{-3})\,.$ (5.10)
This agrees with the spectrum of a theory where each of the $d$ circles is compactified on a circle of radius $N=c^{-1}$.
This achieves the goal of having large dimensions of the torus. This can be applied to each of the three theories. In particular, by regarding the original four dimensional theory as type IIA on $T^{6}$ and considering a black hole that carries large six dimensional momentum charge we can get type IIA in flat ten dimensional space-time. The same theory can be regarded as M-theory on $T^{7}$ and now by switching on appropriate large seven dimensional momentum charge (which corresponds to six dimensional momentum charge and D0-brane charge in the original type IIA description) we can get M-theory in flat eleven dimensional space-time. The same theory can also be regarded as type IIB on $T^{6}$ via a T-duality transformation and by switching on large six dimensional momentum charge (which corresponds to momentum along five of the circles and fundamental string winding along the sixth circle in the type IIA description) we can get type IIB theory in flat ten dimensional space-time.